OpenStudy (anonymous):

Find the derivative. y= (cos2x)^4 - (sin2x)^4

OpenStudy (lgbasallote):

\[y = [\cos (2x)]^4 - [\sin (2x)]^4\] is that the question?

OpenStudy (anonymous):

y'(x) = -4 sin(4 x)

OpenStudy (anonymous):

y'(x) = 2 i e^(4 i x)-2 i e^(-4 i x)

OpenStudy (anonymous):

ans)y'(x) = -8 (sin(2 x) cos^3(2 x)+sin^3(2 x) cos(2 x))

OpenStudy (anonymous):

@lgbasallote is it correct??

OpenStudy (lgbasallote):

nope

OpenStudy (anonymous):

@lgbasallote ,yes

OpenStudy (lgbasallote):

here's a hint.... i'll do the derivative of \[[\cos (2x)]^4\] for you \[\implies 4[\cos (2x)]^3 \times \frac{d}{dx} [\cos (2x)]\] \[\implies 4[\cos (2x)]^3 \times [-\sin (2x)] \times \frac{d}{dx} (2x)\] \[\implies 4[\cos (2x)]^3 \times -\sin (2x) \times 2\] \[\implies -8 [\cos (2x)]^3 \sin (2x)\] does that help?

OpenStudy (anonymous):

@camille08 did u understand???????

OpenStudy (anonymous):

then the derivative of sin2x^4 is 8 (cos 2x)(sin 2x)^3 then we are going to perform the operation.. how

OpenStudy (lgbasallote):

well you can leave it as is

OpenStudy (lgbasallote):

like this: \[\large -8[\cos (2x)]^3 \sin (2x) - 8 \cos (2x) [\sin (2x)]^3\]

OpenStudy (anonymous):

how are we going to combine that

OpenStudy (lgbasallote):

like i said you can leave it like that...

OpenStudy (lgbasallote):

is there an answer you're aiming for it to look like?

OpenStudy (anonymous):

-4sin4x..

OpenStudy (lgbasallote):

factor out.. \[-8 \sin (2x) \cos (2x) [ \cos^2 (2x) + \sin^2 (2x)]\] \[\implies -8\sin (2x) \cos (2x)\] \[\implies -4 [2\sin (2x)\cos (2x)]\] \[\implies -4\sin (2(2x))\] \[\implies -4\sin (4x)\] does that help?

OpenStudy (anonymous):

@lgbasallote , thank you..

OpenStudy (lgbasallote):

welcome :)